OpenStudy (anonymous):

A sofa is on sale for $465, which is 38% less than the regular price. What is the regular price?

OpenStudy (amistre64):

465 = P(1.38) P = 465/1.38 ??

OpenStudy (anonymous):

why 1.38?

OpenStudy (amistre64):

casue I got interest in my head as an error lol

OpenStudy (amistre64):

465 = P - P(.38) = P(1-.38)

OpenStudy (amistre64):

P = 465/(.72)

OpenStudy (anonymous):

For the regular price i got $1,223.68

OpenStudy (anonymous):

$750

OpenStudy (anonymous):

750*.62=465

OpenStudy (amistre64):

645.83 dbl chck: 645.83(.38) = 245.41 645.83 - 245.41 = 400.42

OpenStudy (amistre64):

mines closer :)

OpenStudy (anonymous):

....the sale price is 465...

OpenStudy (amistre64):

right: 465 = P - P(.38) = P(1-.38)

OpenStudy (amistre64):

P = 465/(1-.38)

OpenStudy (amistre64):

.62...my addition was off :)

OpenStudy (amistre64):

750 was the original price